OpenStudy (anonymous):

alegebra 2 help solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

We need to reduce the second equation to less variables. so multiply equation 1 by -4 and add it to equation 2 and tell me what you get.

OpenStudy (anonymous):

8y+9z=39?

OpenStudy (anonymous):

yes, so now we have: -x-y-z = -8 which we can make into x+y+z = 8 by multiplying by -1. 8y+9z =39 2x+2z = 4

OpenStudy (anonymous):

lets get rid of that 2x in equation 3. So multiply my revised equation 1 by -2.

OpenStudy (anonymous):

and add to equation 3

OpenStudy (anonymous):

-2x-2y-2z=-16 -2y=12?

OpenStudy (anonymous):

-2y +2z = -12

OpenStudy (anonymous):

so we have, x+y+z =8 8y+9z =39 -2y+2z = -12

OpenStudy (anonymous):

ok so is that it or is there more steps to it?

OpenStudy (anonymous):

this isn't coming out too nice... is this a basic algebra 2 question?

OpenStudy (anonymous):

yes I couldn't figure it out I got confused

OpenStudy (anonymous):

Well, the answer is x=3 , y=6 , z=−1, but I went through some pretty complex matrixes to find that. I might have missed something though. How were you taught to approach this problem?

OpenStudy (anonymous):

im homeschooled and my mom didn't really understand how to do it either she told me just to put it into a graphing calculator

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

lets start from the beginning again shall we? -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

reducing equation 3 by a factor of 2 gives us: -x-y-z=-8 -4x+4y+5z=7 x+z=2 and we can add equations 1 and 3 together to simply get -x-y-z=-8 -4x+4y+5z=7 -y = -6 y = 6

OpenStudy (anonymous):

now multiply equation 1 by -4 and add it to equation 2:

OpenStudy (anonymous):

4x+4y+4z=32 8y+9z=46

OpenStudy (anonymous):

whoops... it should be 8y+9z=39 now plug the y value we found and solve for z

OpenStudy (anonymous):

The main ideal is that you can eliminate unknown factor from 3 to 2 and then from 2 to 1,you will get the final answer.This is the algebra way,when you are in university,you can use Matrix method to solve the problem. In the above answer,he used equation 1 to eliminate factor x in eqation 2 and 3,then there are only two factors in eqation 2 and 3,you can get y and z,then you put y and z into eqation 1,you will get x.